OpenStudy (anonymous):

x2+x-6/6-3x divided by x2+2x-3/2x+4

Nnesha (nnesha):

ready ?? :)

OpenStudy (anonymous):

yes

Nnesha (nnesha):

\[\huge\rm \frac{ \frac{ x^2 +x-6 }{ 6-3x } }{ \frac{ x^2 +2x -3 }{ 2x+4 } }\] this is question right first it would be much if you change that division the to multiplication

Nnesha (nnesha):

for example \[\large\rm \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } } = \frac{ a }{ b } \times \frac{ d }{ c }\] like this do that first and let me know what you get afte this

OpenStudy (anonymous):

Ok so its x2+x-6/6-3x * 2x+4/x2+2x-3

Nnesha (nnesha):

 yes right now factor quadratic equation at the top left side